Please help me I having trouble with this problem.
Science and medicine.Celsius temperature reading can be converted to Fahrenheit reading using the formula F=9/5 C=32.

Answer by malakumar_kos@yahoo.com(315)


What is the fahrenheit temperature that corresponds to each of the following Celsius temperatures:-10,0,15,100?
The required formula is T in terms of F = (9/5)c+32
When c = 10 F = (9/5)10+32 = 18+32 = 50F
When c = 0 F = (9/5)0+32 = 0+32 = 32F
When c = 15 F = (9/5)15+32 = 27+32 = 59F
When c = 100 F = (9/5)100+32 = 180+32 = 212F